# Anosov diffeomorphisms of open surfaces

### v1

@@ 3.4. Gurevich Entropy is finite
We now begin the proof that $(\widehat{\Sigma},\sigma)$ has finite Gurevich entropy. The reason this is true is essentially that the graph of the associated topological Markov shift has uniformly bounded degree.
@@
$h_{G}(\sigma)\coloneqq\limsup_{n\to\infty}\frac{1}{n}\log\#\{\underline{w}\text{ word of length }n:w_{1}=w_{n}=R\}\in[0,\infty].$
@@ 4. Margulis Measures through Countable Markov Shifts
For an Anosov diffeomorphism of a closed manifold, the measure of maximal entropy is unique and has a number of special properties. One of the most important properties are that its conditional measures along unstable manifolds are conformally invariant under the dynamics and holonomy invariant by stable holonomies. Whereas the conditional of a measure along a foliation is typically not well-defined at every point, the MME gives a system of measures that are well defined along leaves and under local holonomies are invariant on the nose.
@@
In this section we construct a family of Margulis measures, i.e. a family of leaf measures on unstable leaves, which remain invariant up to a fixed multiplicative constant when pushed by the dynamics and which are invariant to holonomies. We wish to construct these measures on $M$ without assuming the existence of a measure of maximal entropy (nor the compactness of $M$). We prove that the Margulis measures are fully supported on unstable leaves, are finite on local unstable leaves, and are infinite on global unstable leaves of periodic points.
The construction of the Margulis measures in this section (Definition 4.12) is general and also applies to codings of non-uniformly hyperbolic systems, such as those studied in [Sar13, BO18, BO20]. Later on we prove some properties for the Margulis measures which make use of the geometry of surfaces.
In this section we work with a one-sided shift spaces, which one may think of as being similar to the space of local unstable manifolds in $\widehat{\Sigma}$.
@@
Conceptually, the right shift on $\widehat{\Sigma}$ is like applying $f^{-1}$ on local unstable leaves. We will also use the following maps.
@@
Let $\tau\colon\widehat{\Sigma}\to\widehat{\Sigma}_{L}$, $\tau((R_{i})_{i\in\mathbb{Z}})=(R_{i})_{i\leq 0}$.
@@
Given a partition set $[R]\subset\widehat{\Sigma}$, for any two points $\underline{R},\underline{S}\in[R]$, we define the local symbolic unstable holonomy map $\widehat{\Gamma}_{\underline{R}\underline{S}}\colon\tau^{-1}[\{\underline{R}\}]\to\tau^{-1}[\{\underline{S}\}]$ by
$\widehat{\Gamma}_{\underline{R}\underline{S}}((\ldots,R_{-1},R_{0})\cdot(x_{0},x_{1},\ldots)):=(\ldots,S_{-1},R_{0})\cdot(x_{0},x_{1},\ldots).$
@@
4.1. A Harmonic Function Depending Explicitly on Smooth Dynamics
In principle, there may be many harmonic functions on $\widehat{\Sigma}$, but not all of them will be easy to relate to the dynamics on $M$. In this subsection we construct a harmonic function, as in Definition 4.5, with explicit values in terms of the dynamics on the manifold. This function will later be used to construct a family of conditional measures.
We will use the same cylinder set notation for the elementary cylinders in $\widehat{\Sigma}_{L}$.
@@ 4.1. A Harmonic Function Depending Explicitly on Smooth Dynamics
For $a\in\mathcal{R}$, define
$P_{n}(a)\coloneqq\left|\mathrm{Per}_{n}\cap[a]\right|,$
where $\mathrm{Per}_{n}$ denotes the set of all periodic chains of (not necessarily minimal) period $n$.
The following properties are clear from the definition of $P_{n}(a)$.
Remark: For any $a\in\mathcal{R}$ and for any $n,m\geq 0$, $P_{n+m}(a)\geq P_{n}(a)P_{m}(a)$; therefore $(\log P_{n}(a))_{n\geq 0}$ is a super-additive sequence.
Definition 4.8.
We say that $\widehat{\Sigma}_{L}$ is recurrent if for some (any) $R\in\mathcal{R}$,
$\sum_{n\geq 0}e^{-nh_{G}(\sigma)}\#\{\underline{w}\text{ word of length }n:w_{1}=w_{n}=R\}=\infty,$
and otherwise we say that it is transient.
Remark: Definition 4.8 is well-defined due to the transitivity of the dynamics on $\widehat{\Sigma}$.
@@
Below the $\pm$ decoration on $\underline{R}^{\pm}$ is meant to emphasize that $\underline{R}^{\pm}$ is a two-sided word.
For $n\geq 1$, set
$W^{s}_{n}(\underline{a})\coloneqq\{\underline{R}^{\pm}\in\widehat{\Sigma}:(R^{\pm}_{k+n})_{k\geq 0}=(a_{k+n})_{k\geq 0}\}.$
For $\underline{R}\in\widehat{\Sigma}_{L}$ we have the local stable sets
$W^{u}(\underline{R}):=\{\underline{R}^{\pm}\in\widehat{\Sigma}:(R^{\pm}_{i})_{i\leq 0}=(R_{i})_{i\leq 0}\}.$
Notice, for every $n\geq 1$, that $\left|W^{u}(\underline{R})\cap W^{s}_{n}(\underline{a}\right|$ is finite and independent of $\underline{R}\in[R_{0}]$. Furthermore, the set $\bigcup_{n\geq 0}W^{u}(\underline{R})\cap W^{s}_{n}(\underline{a})$ is dense in $W^{u}(\underline{R})$.
@@
$Z_{n}(R,\underline{a})\coloneqq\left|W^{u}(\underline{R})\cap W^{s}_{n}(\underline{a})\right|.$
@@
We also define the notation for $R,S\in\mathcal{R}$ we let $Z^{\prime}_{n}(R,S)$ denote the number of valid words of length $n$ of the form $(R,\ldots,S)$.
@@
Fix a periodic point $\underline{a}$. In [Sar01], Sarig shows that for a recurrent, topologically mixing topological Markov shift, there is a sequence $n_{k}$ such that for each $R\in\mathcal{R}$, that
@@ 4.2. Margulis Measures as Conditionals of Generalized Measure of Maximal Entropy
In this section we construct a family of measures on unstable leaves, which are invariant by the dynamics up to a multiplicative constant related to the entropy. When an Anosov diffeomorphism such as ours admits a finite measure of maximal entropy, the conditional measures of the MME can satisfy this property. Even when the system need not admit an MME, one can still define the conditional leaf measures and integrate them in a way which yields an invariant measure that may be infinite or even non-conservative. Hence, the measures we construct can be thought of as the conditionals of a generalized MME.
@@
Let $(\widehat{\Sigma},\sigma_{R})$ be an irreducible, locally compact, topological Markov shift and let $\psi$ an harmonic function that is constant on partition sets. Then there exists a family of Borel probability measures on $\widehat{\Sigma}$, $\{\widehat{p}_{\underline{R}}\}_{\underline{R}\in\widehat{\Sigma}_{L}}$, such that
@@
    For all $\underline{R}\in\widehat{\Sigma}_{L}$, $\widehat{p}_{\underline{R}}$ is carried by, and fully-supported,
@@
    $\widehat{p}_{\underline{R}}\circ\sigma^{-1}=e^{-h_{G}(\sigma)+\log\psi-\log\psi\circ\sigma_{R}}\sum_{\sigma_{R}\underline{S}=\underline{R}}\widehat{p}_{\underline{S}},$
@@
Item (1) and (2) appear explicitly in [BO23, Theorem 5.1] except for the claim of fully support; however this is easily verified for systems that are transitive. Item (3) is a consequence of the following Corollary [BO23, Cor. 5.2].
@@
    For all $\underline{R}\in\widehat{\Sigma}_{L}$, $\widehat{\mu}_{\underline{R}}$ is carried by, and fully-supported on, $\tau^{-1}[\{\underline{R}\}]$,
@@
Definition 4.12 (Margulis measures).
@@
The family $\{\mu_{\underline{R}}\}_{\underline{R}\in\widehat{\Sigma}_{L}}$ is called the Margulis measures.
@@
Note that the above definition only applies to the segment of a local unstable manifold lying within a fixed Markov rectangle $R$; substantial effort will be dedicated to extending holonomy invariance from these holonomies to the true local stable holonomies, i.e. those that do not respect the rectangles.
@@
    For all $\underline{R}\in\widehat{\Sigma}_{L}$, $\mu_{\underline{R}}$ is carried by, and fully-supported on, $W^{u}(x,R_{0})$, for any $x\in\widehat{\pi}[\tau^{-1}[\{\underline{R}\}]]$,
@@
    For all $\underline{R},\underline{S}\in\widehat{\Sigma}_{L}$, $\mu_{\underline{R}}\circ\Gamma_{x,y}^{-1}=\mu_{\underline{S}}$ for any $x\in\widehat{\pi}[\tau^{-1}[\{\underline{R}\}]]$, $y\in\widehat{\pi}[\tau^{-1}[\{\underline{S}\}]]$.
@@
    By the density of $Y$, there exists a point $z\in\mathrm{int}(T)\cap Y$. Consider now the point $w=[x,z]\in W^{u}(x,R)$. For each $i\geq 0$, let $Z_{i}$ be the unique element of $\mathcal{R}$ such that $f^{i}(z)\in\mathrm{int}(Z_{i})\cap Y$. Let $n_{z}\in\mathbb{N}$ be sufficiently large that $f^{-n_{z}}[W^{u}(f^{n_{z}}(w),Z_{n_{z}})]\subseteq T\cap U$. Then,
@@
    The conclusion is immediate when we note that $\Gamma_{x,y}\circ\widehat{\pi}=\widehat{\pi}\circ\widehat{\Gamma}_{\underline{R}\underline{S}}$ on the set $\tau^{-1}([\underline{R}])$.
@@
Note that any element $\underline{R}\in\widehat{\Sigma}_{L}$ is associated to a unique global unstable manifold in $M$: $W^{u}(x)=W^{u}(y)$ for any $x,y\in\hat{\pi}(\tau^{-1}(\underline{R}))$. We denote this unstable manifold by $W^{u}(\underline{R})$. In a sense, this unstable manifold is the image of a combinatorial unstable manifold. Hence one would like to define the Margulis measures on $W^{u}(\underline{R})$ by pushing forward the measure from the combinatorial unstable manifold to $M$. In order to do this, one must at least know that the resulting measure is locally finite. This is guaranteed by the following claim.
@@
By Lemma 3.20, there exists a constant $C$ such that any point $x\in M$ has at most $C$ pre-images under $\widehat{\pi}$. There exist a finite number of local leaves $W^{u}(x,T_{i})$ that $W^{u}(x,R)$ intersects. It then suffices to show that for each $T_{i}$ that $\mathcal{P}(W^{u}(x,T_{i}))$ is finite. Suppose that there there exist distinct codings $\underline{S}^{1},\ldots,\underline{S}^{n}\in\widehat{\Sigma}_{L}\in\mathcal{P}(W^{u}(x,T_{i}))$, where $n>C$. Let $\underline{R}$ be a coding of $x$ with $R_{0}=T_{i}$ and $\underline{R}^{\prime}$ be $(R_{i})_{i\geq 0}$ then by the Markov property for each $1\leq i\leq n$, the word $\underline{S}^{i}\cdot\underline{R}^{\prime}\in\widehat{\Sigma}$ is valid. Moreover, each of these words codes the point $x$ as they code the unique point in $W^{s}(x,T_{i})\cap W^{u}(\underline{S})$. Thus we have found $n>C$ distinct words coding the point $x$; a contradiction. The result follows. ∎
@@
is well-defined as an increasing sum of measures, and is finite on any finite arc $I\subseteq W^{u}(\underline{R})$ where $W^{u}(\underline{R})$ is defined above.
@@
The recurrent part of $\widehat{\Sigma}_{L}$, which we denote $\widehat{\Sigma}_{L}^{\#}$ is the set of words in $\widehat{\Sigma}_{L}$ that contain some particular symbol infinitely many times.
In particular, the recurrent part of $\widehat{\Sigma}_{L}$ contains all the periodic chains in $\widehat{\Sigma}_{L}$. Note that $\tau^{-1}[\widehat{\Sigma}_{L}^{\#}]$ is exactly the set of all codings of all points which return to some compact subset of $M$ infinitely often under the forward iteration of $f^{-1}$.
@@
Let $\underline{R}\in\widehat{\Sigma}_{L}^{\#}$, and let $n_{k}$ be an increasing sequence such that $\sigma_{R}^{n_{k}}\underline{R}\in[R_{n_{0}}]$ for all $k\geq 0$. Then,

### v2

@@ 2.4. Gurevich Entropy is finite
We now begin the proof that $(\widehat{\Sigma},\sigma)$ has finite Gurevich entropy. The Gurevich entropy is a generalization of topological entropy to the non-compact setting, which measures the growth rate of the number of periodic orbits of length $n$ that intersect a small neighborhood. In the classical Anosov case, one can similarly define entropy by studying returns to a small open set. One can check its equivalence to the usual definition using Bowen balls by using the shadowing property. In the case of a countable Markov shift, as long as the dynamics are transitive, the Gurevich entropy is independent of the neighborhood used. (For a detailed discussion see [Sar09, §3.1.3]). The reason that the Gurevich entropy is finite in this case is essentially that the graph of the associated topological Markov shift has uniformly bounded degree.
@@
$h_{G}(\sigma)\coloneqq\limsup_{n\to\infty}\frac{1}{n}\log\#\{\underline{w}\text{ word of length }n:w_{0}=w_{n-1}=R\}\in[0,\infty].$
@@ 3. Margulis Measures through Countable Markov Shifts
For an Anosov diffeomorphism of a closed manifold, the measure of maximal entropy is unique and has a number of special properties. Some of the most important properties are that its conditional measures along unstable manifolds are conformally invariant under the dynamics and holonomy invariant by stable holonomies. See, for example, the discussion in [KH97]. Whereas the conditional of a measure along a foliation is typically not well-defined at every point, the MME gives a system of measures that are well defined along leaves and under local holonomies are invariant.
@@
In this section we construct a family of Margulis measures, i.e. a family of leaf measures on unstable leaves, which remain invariant up to a fixed multiplicative constant when pushed by the dynamics and which are invariant under holonomies. We wish to construct these measures on $M$ without assuming the existence of a measure of maximal entropy (nor the compactness of $M$). We prove that the Margulis measures are fully supported on unstable leaves, are finite on local unstable leaves, and are infinite on global unstable leaves of periodic points.
The construction of the Margulis measures in this section (Definition 3.11) is general and also applies to codings of non-uniformly hyperbolic systems, such as those studied in [Sar13, BO18, BO20].
3.1. Coding and Associated One Sided Shifts
We will fix for the rest of the section a Markov partition $\mathcal{R}$ and associated coding $\hat{\pi}\colon\widehat{\Sigma}\to M$ as in Theorem 2.20. In our construction, we will also work with two kinds of local-unstable sets, stable sets in $\widehat{\Sigma}$ as well as local unstable manifolds in $M$. The dynamics on the space of local unstable sets in $\widehat{\Sigma}$ is given by the dynamics on an associated $1$-sided shift, which we now describe.
@@ 3.1. Coding and Associated One Sided Shifts
Conceptually, the right shift on $\widehat{\Sigma}_{L}$ is like applying $f^{-1}$ on the set of local unstable leaves. We will also use the following maps, which relate $\widehat{\Sigma}$ and $\widehat{\Sigma}_{L}$.
@@
Let $\tau\colon\widehat{\Sigma}\to\widehat{\Sigma}_{L}$ be the map $\tau((R_{i})_{i\in\mathbb{Z}})=(R_{i})_{i\leq 0}$. Note that for a word $\underline{R}\in\widehat{\Sigma}_{L}$, that $\tau^{-1}[\{\underline{R}\}]$ is a (symbolic) local unstable set in $\widehat{\Sigma}$. Hence we will also denote this set by $W^{u}_{0}(\underline{R})$.
Next we will have the symbolic stable holonomies between local unstable sets.
@@
Given a cylinder set $[R]\subset\widehat{\Sigma}$, for any two points $\underline{R},\underline{S}\in[R]\subset\widehat{\Sigma}_{L}$, we define the local symbolic stable holonomy map $\widehat{\Gamma}_{\underline{R}\underline{S}}\colon W^{u}_{0}(\underline{R})\to W^{u}_{0}(\underline{S})$ by
$\widehat{\Gamma}_{\underline{R}\underline{S}}((\ldots,R_{-1},R_{0})\cdot(x_{0},x_{1},\ldots))\coloneqq(\ldots,S_{-1},R_{0})\cdot(x_{0},x_{1},\ldots).$
@@
3.2. A Harmonic Function on the topological Markov shift
In the classical approach to the construction of the MME for an Anosov diffeomorphism of a compact manifold, one first constructs a harmonic function of a particular Ruelle operator by exhibiting that the operator has spectral gap. Similarly, we will construct the Margulis measures by using a harmonic function for the Ruelle operator. However, in the non-compact case when the topological Markov shift is transient the argument from the compact case has a number of complications. A priori there is no obvious reason why any harmonic functions should exist. For example, in the non-compact case the Ruelle operator may not have a spectral gap nor need it be exponentially mixing. Hence to obtain a harmonic function more delicate work is required. Fortunately, work of Cyr [Cyr10] and Shwartz [Shw19] shows that such harmonic functions do exist, and in the transient case there may be many of them. There is then another potential issue: the structure of the harmonic function might have little relation to the structure of the topological Markov shift and hence the associated dynamics on the coded space. This might cause problems when we attempt to turn information about holonomy invariance in the shift space into holonomy invariance on the manifold $M$. Fortunately, the way that Cyr constructs a harmonic function by counting intersections of stable sets in the topological Markov shift has a suitable dynamical structure. Later, in Claim 3.29, we will use this specific form of the harmonic functions to verify the holonomy invariance of the Margulis measures by relating these intersections in $\widehat{\Sigma}$ to the intersections of a curve in $M$ with a long stable manifold.
Let us now contrast this with a more classical approach to the construction of Margulis measures such as Hiraide considers [Hir01]. To construct the Margulis measures along a piece of unstable manifold $W^{u}_{loc}(x)$, Hiraide takes a piece of stable manifold $W^{s}_{L}(y)$ of length $L$ and places a $\delta$-mass at each intersection point of $W^{s}_{L}(y)$ with $W^{u}_{loc}(x)$. Then Hiraide normalizes and takes a weak* limit. Broadly, this is similar to the construction of Cyr. However, in the non-compact setting, unless one chooses the leaf $W^{s}_{L}(y)$ carefully, problems may arise. For example, in our setting it is possible that $W^{s}(y)$ might intersect $W^{u}_{loc}(x)$ only finitely many times, which would result in a fully supported measure.
We now proceed with the construction of the harmonic functions. We will use the same cylinder set notation for the elementary cylinders in $\widehat{\Sigma}_{L}$.
@@ 3.2. A Harmonic Function on the topological Markov shift
We say that $\widehat{\Sigma}_{L}$ is recurrent if for some $R\in\mathcal{R}$,
(2) $\sum_{n\geq 0}e^{-nh_{G}(\sigma)}\#\{\underline{w}\text{ word of length }n:w_{0}=w_{n-1}=R\}=\infty,$
and otherwise we say that it is transient. Definition 3.7 is well-defined due to the transitivity of the dynamics on $\widehat{\Sigma}$, hence it is equivalent to say that $\hat{\Sigma}$ is recurrent if (2) holds for all $R\in\mathcal{R}$.
@@
We will use both local stable and local unstable sets in $\hat{\Sigma}$ and count intersections between them. We already introduced the symbolic local unstable sets in Definition 3.2, so we now introduce the local stable sets. For $n\geq 0$, and $\underline{a}\in\widehat{\Sigma}$, we have local stable sets of different diameters. Set
$W^{s}_{n}(\underline{a})\coloneqq\{\underline{R}^{\pm}\in\widehat{\Sigma}:(R^{\pm}_{k+n})_{k\geq 0}=(a_{k+n})_{k\geq 0}\}=\sigma^{-n}(W^{s}_{0}(\underline{a})).$
Notice, for every $n\geq 1$, that $\left|W^{u}_{0}(\underline{R})\cap W^{s}_{n}(\underline{a})\right|$ is finite and independent of $\underline{R}\in[R_{0}]\subset\widehat{\Sigma}_{L}$. Furthermore, the set $\bigcup_{n\geq 0}W^{u}_{0}(\underline{R})\cap W^{s}_{n}(\underline{a})$ is dense in $W^{u}_{0}(\underline{R})$.
@@
$Z_{n}(R,\underline{a})\coloneqq\left|W^{u}_{0}(\underline{R})\cap W^{s}_{n}(\underline{a})\right|.$
@@
We also define the notation for $R,S\in\mathcal{R}$. We let $Z^{\prime}_{n}(R,S)$ denote the number of valid words of length $n$ of the form $(R,\ldots,S)$.
@@
Fix a periodic point $\underline{a}$. In [Sar01], Sarig shows that for a recurrent, topologically mixing topological Markov shift, there is a sequence $n_{k}\to+\infty$ such that for each $R\in\mathcal{R}$, that
@@ 3.3. Margulis Measures as Conditionals of Generalized Measure of Maximal Entropy
In this section we construct a family of measures on unstable leaves that are invariant by the dynamics up to a multiplicative constant related to the entropy. When an Anosov diffeomorphism such as ours admits a finite measure of maximal entropy, the conditional measures of the MME can satisfy this property. Even when the system need not admit an MME, one can still define the conditional leaf measures and integrate them in a way which yields an invariant measure that may be infinite or even non-conservative. Hence, the measures we construct can be thought of as the conditionals of a generalized MME.
@@
Let $(\widehat{\Sigma},\sigma_{R})$ be an irreducible, locally compact, topological Markov shift, $(\widehat{\Sigma}_{L},\sigma_{R})$ the associated $1$-sided shift, and let $\psi$ be a harmonic function that is constant on partition sets. Then there exists a family of Borel probability measures on $\widehat{\Sigma}$, $\{\widehat{p}_{\underline{R}}\}_{\underline{R}\in\widehat{\Sigma}_{L}}$, such that
@@
    For all $\underline{R}\in\widehat{\Sigma}_{L}$, $\widehat{p}_{\underline{R}}$ is carried by, and fully-supported, on $\tau^{-1}[\{\underline{R}\}]=W^{u}_{0}(\underline{R})$.
@@
    $\widehat{p}_{\underline{R}}\circ\sigma^{-1}=\sum_{\sigma_{R}\underline{S}=\underline{R}}e^{-h_{G}(\sigma)+\log\psi(\underline{S})-\log\psi\circ\sigma_{R}(\underline{S})}\widehat{p}_{\underline{S}},$
@@
Items (1) and (2) appear explicitly in [BO23, Theorem 5.1], where the full support property is an immediate consequence of item (2): $\widehat{p}_{\underline{R}}([R_{0},a_{1},\ldots,a_{n-1}])=e^{-nh_{G}(\sigma)}\frac{\psi(R_{0})}{\psi(a_{n-1})}>0$. Item (3) is a consequence of the following Corollary [BO23, Corollary 5.2].
@@
    For all $\underline{R}\in\widehat{\Sigma}_{L}$, $\widehat{\mu}_{\underline{R}}$ is carried by, and fully-supported on, $\tau^{-1}[\{\underline{R}\}]=W^{u}_{0}(\underline{R})$,
@@
Definition 3.11 (Local Margulis measures).
@@
The family $\{\mu_{\underline{R}}\}_{\underline{R}\in\widehat{\Sigma}_{L}}$ is called the local Margulis measures.
@@
Note that the above definition only applies to the segment of a local unstable manifold lying within a fixed Markov rectangle $R$; substantial effort will be dedicated to extending holonomy invariance from these holonomies to the full stable holonomies on $M$, i.e. those that do not respect the rectangles.
@@
    For all $\underline{R}\in\widehat{\Sigma}_{L}$, $\mu_{\underline{R}}$ is carried by, and fully-supported on, $W^{u}(x,R_{0})$, for any $x\in\widehat{\pi}[W^{u}_{0}(\underline{R})]$,
@@
    For all $\underline{R},\underline{S}\in\widehat{\Sigma}_{L}$, $\mu_{\underline{R}}\circ\Gamma_{x,y}^{-1}=\mu_{\underline{S}}$ for any $x\in\widehat{\pi}[W^{u}_{0}(\underline{R})]$, $y\in\widehat{\pi}[W^{u}_{0}(\underline{S})]$.
@@
    Let $Y$ be as in Definition 2.14. By the density of $Y$, there exists a point $z\in\mathrm{int}(T)\cap Y$. Consider now the point $w=[x,z]\in W^{u}(x,R)$. For each $i\geq 0$, let $Z_{i}$ be the unique element of $\mathcal{R}$ such that $f^{i}(z)\in\mathrm{int}(Z_{i})\cap Y$. Let $n_{z}\in\mathbb{N}$ be sufficiently large that $f^{-n_{z}}[W^{u}(f^{n_{z}}(w),Z_{n_{z}})]\subseteq T\cap U$. Then,
@@
    The conclusion is immediate when we note that the symbolic and topological holonomies are intertwined by $\widehat{\pi}$, i.e.  $\Gamma_{x,y}\circ\widehat{\pi}=\widehat{\pi}\circ\widehat{\Gamma}_{\underline{R}\underline{S}}$ on the set $W^{u}_{0}(\underline{R})$.
@@
Note that any element $\underline{R}\in\widehat{\Sigma}_{L}$ is associated to a unique global unstable manifold in $M$: $W^{u}(x)=W^{u}(y)$ for any $x,y\in\hat{\pi}(W^{u}_{0}(\underline{R}))$. We denote this unstable manifold by $W^{u}(\underline{R})$. (Note that this is distinguished from the combinatorial local stable manifold by the lack of the 00 subscript). Define $W^{s}(\underline{R})$ analogously for $\underline{R}\in\widehat{\Sigma}_{R}$. This unstable manifold is the image of a symbolic unstable manifold. Hence one would like to define the Margulis measures on $W^{u}(\underline{R})$ by pushing forward the measure from the combinatorial unstable manifold to $M$. In order to do this, one must at least know that the resulting measure is locally finite. This is guaranteed by the following claim.
@@
By Lemma 2.19, there exists a constant $C$ such that any point $x\in M$ has at most $C$ pre-images under $\widehat{\pi}$. There exist a finite number of local leaves $W^{u}(x,T_{i})$ that $W^{u}(x,R)$ intersects. It then suffices to show that for each $T_{i}$ that $\mathcal{P}(W^{u}(x,T_{i}))$ is finite. Suppose that there there exist distinct codings $\underline{S}^{1},\ldots,\underline{S}^{n}\in\widehat{\Sigma}_{L}\in\mathcal{P}(W^{u}(x,T_{i}))$, where $n>C$. Let $\underline{R}$ be a coding of $x$ with $R_{0}=T_{i}$ and $\underline{R}^{\prime}$ be $(R_{i})_{i\geq 0}$ then by the Markov property, for each $1\leq i\leq n$, the word $\underline{S}^{i}\cdot\underline{R}^{\prime}\in\widehat{\Sigma}$ is valid. Moreover, each of these words codes the point $x$ as they code the unique point in $W^{s}(x,T_{i})\cap W^{u}(\underline{S})$. Thus we have found $n>C$ distinct words coding the point $x$; a contradiction. The result follows. ∎
@@
is well-defined as an increasing sum of measures and is finite on any finite arc $I\subseteq W^{u}(\underline{R})$ where $W^{u}(\underline{R})$ is defined above to be the full unstable manifold of $\widehat{\pi}(\underline{R})$.
@@
Later we will need the following recurrence properties of $\widehat{\Sigma}_{L}$.
@@
The recurrent part of $\widehat{\Sigma}_{L}$, which we denote $\widehat{\Sigma}_{L}^{\#}$, is the set of words in $\widehat{\Sigma}_{L}$ that contain some particular symbol infinitely many times.
In particular, the recurrent part of $\widehat{\Sigma}_{L}$ contains all the periodic chains in $\widehat{\Sigma}_{L}$. Note that $\tau^{-1}[\widehat{\Sigma}_{L}^{\#}]$ is exactly the set of all codings of all points that return to some compact subset of $M$ infinitely often under the forward iteration of $f^{-1}$.
@@
[…]